OpenStudy (anonymous):

What values for q (0 < q < 2p) satisfy the equation? 2sqrt(2)sin q + 2 = 0

OpenStudy (jdoe0001):

\(\bf \large 2\sqrt{2}sin(q)+2=0\ \ ;\ \ 0 < q < 2\pi\)

OpenStudy (anonymous):

Can someone teach me how to solve this? I would like to be able to solve it on my own bu the text book is just not making sense today.

OpenStudy (anonymous):

do the regular algebra by isolating sin(q) on one side of the equal sign....

OpenStudy (jdoe0001):

\(\bf \large {2\sqrt{2}sin(q)+2=0; 0 < q < 2\pi\\ sin(q) = \frac{2}{2\sqrt{2}} = \frac{1}{\sqrt{2}} = \frac{\sqrt{2}}{2}}\)

OpenStudy (anonymous):

that should be negative on the right side... ie, \(\frac{-2}{2\sqrt2} \)

OpenStudy (anonymous):

wait! so the first step is to divide by \[2\sqrt{2}\] to move it to the right??

OpenStudy (jdoe0001):

actually, yes, should be -2 :/

OpenStudy (anonymous):

no... move the lone "2" first by subtractig...

OpenStudy (anonymous):

ok i see that which is where the nominator comes from.

OpenStudy (jdoe0001):

\(\bf \large 2\sqrt{2}sin(q)+2=0; 0 < q < 2\pi\\ \large sin(q) = -\frac{2}{2\sqrt{2}} = -\frac{1}{\sqrt{2}} = -\frac{\sqrt{2}}{2}\)

OpenStudy (anonymous):

yep...

OpenStudy (anonymous):

how did the \[\sqrt{2}\] get on top of the fraction?

OpenStudy (jdoe0001):

then for that, just check your Unit Circle for the angles between 0 and \(\bf \large \pi\) that have that value with a NEGATIVE sign

OpenStudy (jdoe0001):

it gets there by "simplifying" the rational, so-called

OpenStudy (jdoe0001):

\(\bf \large { 2\sqrt{2}sin(q)+2=0; 0 < q < 2\pi\\ sin(q) = -\frac{2}{2\sqrt{2}} = -\frac{1}{\sqrt{2}} \times \frac{\sqrt{2}}{\sqrt{2}} = -\frac{\sqrt{2}}{2} } \)

OpenStudy (anonymous):

Alright, none of my choices are negative though.

OpenStudy (jdoe0001):

well, check your Unit Circle for such values

OpenStudy (jdoe0001):

between 0 and \(\bf \large \pi \) there'd be 2 angles with that sine value

OpenStudy (anonymous):

I'm sorry I have got to sound dumb but I don't know what a unit circle is.

OpenStudy (anonymous):

All of my options have two fractions. I am not sure how to paste the pictures to here.

OpenStudy (jdoe0001):

http://i.stack.imgur.com/r8uHr.gif <---

OpenStudy (jdoe0001):

many Unit Circles online, you'd need one :)

OpenStudy (anonymous):

Thank you! Now you said between 0 and pi there would be angles with the sign value.

OpenStudy (jdoe0001):

yes, and at those angles, is where the equation holds, that is \(\bf 2\sqrt{2}sin(q)+2=0\)

OpenStudy (anonymous):

So my answer would be \[2\sqrt{2}\]

OpenStudy (anonymous):

oops i meant to flip those

OpenStudy (anonymous):

\[\sqrt{2}/2\]

OpenStudy (jdoe0001):

well, "q" is an angle unit, so the values where \(\bf \large sin(q) = -\frac{\sqrt{2}}{2}\) is what "q" equals to, there are 2 angles, one in the 3rd quadrant and another in the 4th quadrant

OpenStudy (anonymous):

ok so one of my answers should be \[7\pi/4\]

OpenStudy (anonymous):

and the other will be found at 135deg \[\pi/4\]

OpenStudy (anonymous):

!I see it. Thank you so much I actually understand how to do these now!!!